A hot water radiator at 310k temperature radiates thermal radiates like a black body. Its total surface area is 1.6 m^2. Find the thermal power radiated by it.
1
Expert's answer
2020-09-03T07:14:45-0400

Stefan’s Law of Radiation (black body):

"E = \u03c3AT^4"

A is the surface area, m2

T is the surface temperature of the black body, K

Stefan’s constant "\sigma = 5.670 \times 10^{−8} \frac {W}{( m^2 ⋅ K^4 )}"

"A = 1.6 m^2"

"T = 310 K"

"E = 5.670 \times 10^{−8}\times1.6\times 310^4 = 5.670 \times 10^{−8}\times 1.6\times 9.235 \times 10^{9} = 837.8 W"

Answer: 837.8 W